OpenStudy (anonymous):

The volume (V) of a cylinder is V = π r2 h. Find h, if V = 141.3, π = 3.14, and r = 3. Round to the nearest tenth if necessary.

OpenStudy (anonymous):

u gotta be kidding... just plugging those numbers into ur calculator..

OpenStudy (anonymous):

what kind of question is that .... u alrdy have the Volume in ur question..

OpenStudy (anonymous):

you lost me already...

OpenStudy (anonymous):

u alrdy have all the information in the question

OpenStudy (anonymous):

h = V divided by pi r^2

OpenStudy (anonymous):

now plug in those numbers and tadaaa~~ u got ur answer

OpenStudy (theviper):

\(\Large{V=\pi r^2h}\) \(\huge{h=\frac{V}{\pi r^2}}\)

OpenStudy (anonymous):

<3 ohh yea thanks sorry takin a 120 question online course and startin to get mind bottled

OpenStudy (theviper):

no problem ;)